OpenStudy (anonymous):

what is the general solution of (xy^3)dx + (e^(x^2))dy = 0 my progress: (xy^3)dx = -(e^(x^2))dy xdx/(e^(x^2)) = dy/y^3 ...... i know how to integrate dy/y^3 but does anyone know how to integrate xdx/(e^(x^2)) ?

OpenStudy (shubhamsrg):

let x^2 = t see if this helps

OpenStudy (anonymous):

uh...lemme try

OpenStudy (anonymous):

It's quite easy if you just get the hang of how to do it.

OpenStudy (anonymous):

wait, does that mean -dx/(e^(x^2)) becomes -t/(e^(t^2)) ?

OpenStudy (shubhamsrg):

nop, try again

OpenStudy (anonymous):

No, keep trying.

OpenStudy (anonymous):

so if x^2 = t then -dx/(e^(x^2)) becomes -dx/(e^t) right? then I integrated via ln therefore -dx/(e^t) becomes lne^t then ln(e^(x^2))?

OpenStudy (anonymous):

uhm...is it right?

OpenStudy (shubhamsrg):

nop it becomes dt/2e^t please recheck

OpenStudy (anonymous):

since the integration of (e^(x^2)) [where e^u du = e^u + C] is 2(e^(x^2)) right? i was kinda thinking of that as well but what I was wondering was how to integrate them fully. because 2(e^(x^2) ) is not the derivative of dx [meaning dx/x = ln|x| +C ] so what should I do with (e^(x^2)) so it can be a derivative of dx?

OpenStudy (anonymous):

wait, I think i'm getting it. since -dx/(e^(x^2)) becomes dt/2e^t then it becomes (-1/2)(dt/e^t) right? then: (-1/2)(e^-t)(dt) = (1/2)(e^-t) + C right?

OpenStudy (shubhamsrg):

integration of (e^(x^2)) [where e^u du = e^u + C] is NOT 2(e^(x^2)).. and yes, latter thing you said was right

OpenStudy (anonymous):

thank you! (oh and (e^(x^2)) = 2x(e^(x^2)) right? sorry. my bad ^^;;)

OpenStudy (phi):

the derivative of \[ d\ e^{-x^2} = e^{-x^2} d\ (-x^2) = -e^{x^2} 2x \ dx\]